OpenStudy (anonymous):

one roofer can put a new roof on a house three times faster than another. working together, they can roof a house in 4 days. how long would it take the faster roofer working alone? it seems so simple but for the life of me i cant figure it out, please help

OpenStudy (anonymous):

roofer A is 3 times fatster than roofer B A = 3B A + B = a roof in 4 days substitute B for 3A A + 3A = 4 days 4 A = 4 days it woud take 4 times as long, so 16 days? im not sure at all though!

OpenStudy (anonymous):

no israrani, that is incorrect.

OpenStudy (anonymous):

it should be a+a/3 in 4 days, not a + 3a in 4 days

OpenStudy (anonymous):

oh my bad! just found this site on stumbleupon and thought ild try one really quick haha sorry!

OpenStudy (anonymous):

sophie, are you still interested in the answer or no?

OpenStudy (anonymous):

one roofer can put a new roof on a house three times faster than another. Working together, they can roof a house in 4 days. How long would it take the faster roofer working alone? Let x be the fastest roofers time to roof the house Then the slower roofers time will be 3x Use the following formula to calculate how the work is done: \[\frac{1}{a}+\frac{1}{b}=\frac{1}{c}\] where A & B are the times done by themselves, and C is the time it takes them to do the work together, thus we have: \[\frac{1}{x}+\frac{1}{3x}+\frac{1}{4}\]\[\frac{4}{3x}=\frac{1}{4}\]\[3x=16\]\[x=\frac{16}{3}\]So the fast roofer can do it by himself in about 5.33 days!

OpenStudy (anonymous):

\[100/4 = 25\] How much roof is laid in one day (%) \[25/4 = 6.25\] How much roof is laid by the useless roofer (%) \[6.25 * 3=18.75\] How much roof is laid by the godlike roofer (%) \[100/18.75= \] how many days (roughly 5.33)

OpenStudy (anonymous):

yes definitly intrested in the answer

OpenStudy (anonymous):

well, sophie, nancy has given the answer in a nice form.

OpenStudy (anonymous):

yes she has..thank you very much!!

OpenStudy (anonymous):

click good answer if u like my answer,ty